Let a1, a2, a3....be terms of an AP. If (a1 +a2.....ap)/(a1 +a2.....aq) =p^2/q^2. ,p is not equal to q. Then a6/a21 must be. A) 2/7 B) 11/41 C) 7/2

Grade:11

2 Answers

SREEKANTH
7 years ago
(a1 +a2.....ap)/(a1 +a2.....aq) =p^2/q^2
p/2(a1+(p-1)d)/q/2(a1+(q-1)d)=p^2/q^2
(a1+(p-1)d)/(a1+(q-1)d)=p/q
q(a1+(p-1)d)=p(a1+(q-1)d)
(p-q)a1=(q-p)d
a1=-d
a6/a21=(a1+5d)/(a1+20d)
           =-4a1/-19a1
           =4/19 is the answer
Vanshika agrawal
6 years ago
(a1 +a2.....ap)/(a1 +a2.....aq) =p^2/q^2p/2(2a1+(p-1)d)/q/2(2a1+(q-1)d)=p^2/q^2(2a1+(p-1)d)/(2a1+(q-1)d)=p/qq(2a1+(p-1)d)=p(2a1+(q-1)d)(p-q)2a1=(p-q)d2a1=da6/a21=(a1+5d)/(a1+20d) =a+10a/a+40a =11/41 is the answer